A number of years after Andile deposited the R1000 into the second account, the accrued amount in the second account is 1605.What is the time the money was invested for in years?

1
Expert's answer
2022-03-01T12:24:20-0500

I = PRT/100

I = amount - principal

I= R1605 - R1000

=605.

Therefore, 605 = PRT/100

(605*100) = (1000*11)T

T = 60500/11000

= 5.50 years